OpenStudy (anonymous):

Help? Which property is used in Step 1 of simplifying the expression 5(z-x)+z+0? Expression: 5(z – x) + z + 0 Step 1: 5z – 5x + z + 0 Step 2: 5z + z – 5x + 0 Step 3: 5z + 1z – 5x + 0 Step 4: 6z – 5x + 0 Step 5: 6z – 5x A Zero Property of Multiplication B Identity Property of Addition C Distributive Property D Associative Property of Multiplication (Thanks)

OpenStudy (anonymous):

Here are examples of each. Choose which one. A. Zero Property of Multiplication 0 * x = 0 B. Identity Property of Addition 0 + 5 = 5 C. Distributive Property 3(x + 5) = 3x + 15 D. Associative Property of Multiplication (3 * 2) * 5 = 3 * (2 * 5) Just in case you're wondering, * means multiplication.